Given a cylinderical tin with radius of each cross section 3m and hight 6m. Calculate it's Surface Area.
a. The two (2) ends are not clossed.
b. One (1) end is closed.
c. The two (2) ends are closed.

Answer by stanbon(75887)

Given a cylinderical tin with radius of each cross section 3m and hight 6m. Calculate it's Surface Area.
a. The two (2) ends are not closed.
Ans:: S = 2*pi*r*h = 2pi*3*6 = 36pi sq. meters
------------------------------------
b. One (1) end is closed.
Ans:: S + pi*3^2 = 36pi+ 9pi = 45pi sq meters
-------------------------------
c. The two (2) ends are closed.
Ans:: S + 2(pi)*9 = 45+9pi = 54pi sq meters
-------------
Cheers,
Stan H.
------------------